OpenStudy (anonymous):

If the unicorn seat on a merry–go–round is 5 feet from the center, and the merry–go–round completes a full rotation in 6 seconds, how many feet does the unicorn travel in one second a. 5π/4 feet b.π/6 radians c.13.08 feet d.5.23 feet

OpenStudy (anonymous):

what do you think it is?

OpenStudy (anonymous):

or do you want me to work through it?

OpenStudy (anonymous):

5 feet from the center means that your radius is equal to 5. r = 5 Time is 6 seconds. t = 6 First you need to find the total distance the unicorn travels in 1 rotation, which in this case is the circumference of a circle with a radius of 5.

OpenStudy (anonymous):

C = 2*(pi)*r C = 10(pi) So the unicorn travels 10(pi) feet in 6 seconds. Now just have to divide. 10(pi)/6 In one second he travels, (5(pi))/2 feet, or about_________

OpenStudy (anonymous):

5.24 feet.

OpenStudy (anonymous):

Thank you. sorry i got off for a minute